# Self-Supervised Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography

### v3

@@ 4.2 Leaf Phantom
We use the phantom data from [6], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left of Fig. 5a and details in Sec. S1.2). The reconstructions of our methods and the baselines are shown in Fig. 5a. The lower left of the phantom is dim due to nonuniform illumination [6]. We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS. Ours are able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak and DAS and APACT both fail. The SOS reconstructed by our methods have a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole. PG still suffers from stripe-like artifacts in the SOS (zoom in) but also reconstructs a well-resolved IP image.
@@ 4.4 Ablation Studies
Our ablation studies focus on the effects of the number of delays $M$, TV regularization $\lambda$, and network size. We observe an increase in the performance at the cost of compute time as more extra delay distances are used and more redundancy is created in the DAS stack. This improvement gradually diminishes beyond a delay count of 16, which we selected for our implementation (see Fig. 6). For NF network size (see Fig. S3 and Sec. S3.3), our method fails either when the network is too small (NF’s implicit smoothness regularization is too strong) or large (the NF has a stronger ability to learn high-frequency features and has weaker implicit regularization). For the TV regularization of PG (see Fig. S2 and Sec. S3.2), a larger $\lambda$ encourages smoother SOS results and reaches optimal performance at a weight of $1\times 10^{-4}$.
@@ 5 Discussion and Conclusion
We have introduced a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and speed of sound in PACT. This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multichannel image representation analogous to a focal stack. We highlight the following as the key features of our novel method:
@@
Generalizability. Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems [4]. In 3D PACT, where multiple 2D scans are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which drastically scales up the computational time and memory requirements. To address this challenge, our efficient NF SOS representation combined with efficient and differentiable forward model offers a promising solution. By reducing the degrees of freedom and lowering computational costs, the proposed method could enable practical 3D PACT applications.
@@
S1.3 In vivo Mouse Liver
The PA signals in our in vivo experiments were collected from a nude mouse liver at 1064 nm [3]. The water temperature was 31C during the experiment, thus we used a background SOS of 1511.4 m/s. See figure S1 to see the alignment between recovered SOS and IP, which matches better in our method than others.
S1.4 Image Stitching
@@
One critical contribution of our work is the improved wavefront model compared to the previous work APACT [6]. As explained in the main text, the wavefront model represents the wavefront advancement or retardation induced by a nonuniform speed‐of‐sound distribution in tissue relative to an ideal spherical wave in a homogeneous medium. Mathematically, the wavefront function $w(\theta)$, which depicts the propagation length difference relative to the spherical wave when received by the transducers, is a function of the propagation direction $\theta$. This wavefront function can be expanded using a Fourier series:
@@
In Fig. S2, we show the effects of TV regularization weight $\lambda$ with our pixel grid SOS representation. We observe that the IP image quality drops when $\lambda>1\times 10^{-4}$, while the SOS accuracy reaches optimum around $\lambda=1\times 10^{-4}$. Therefore, we select $\lambda>1\times 10^{-4}$ (marked with a green star in Fig. S2) for all reconstructions in our implementation of PG.
@@
We provide six videos showing the convergence process of the IP image and SOS for our methods with our numerical phantom (nf_numerical_phantom.mp4, and pg_numerical_phantom.mp4), leaf phantom (nf_leaf_phantom.mp4, and pg_leaf_phantom.mp4), and in vivo mouse liver (nf_in_vivo_mouse_liver.mp4 and pg_in_vivo_mouse_liver.mp4).
Figure S4: Results on all five numerical phantoms.

### v4

@@ 4.2 Leaf Phantom
We use the phantom data from [5], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left of Fig. 5a and details in Sec. S1.2). We show the EIR and MTF of our ring array system in and Sec. S1.4. The reconstructions of our methods and the baselines are shown in Fig. 5a. The lower left of the phantom is dim due to nonuniform illumination [5]. We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS. Ours are able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak and DAS and APACT both fail. The SOS reconstructed by our methods have a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole. PG still suffers from stripe-like artifacts in the SOS (zoom in) but also reconstructs a well-resolved IP image.
@@ 4.4 Ablation Studies
Our ablation studies focus on the effects of the number of delays $M$, TV regularization $\lambda$, and network size. We observe an increase in the performance at the cost of compute time as more extra delay distances are used and more redundancy is created in the DAS stack. This improvement gradually diminishes beyond a delay count of 16, which we selected for our implementation (see Fig. 6). For NF network size (see Fig. S4 and Sec. S3.3 in the supplement), our method fails either when the network is too small (NF’s implicit smoothness regularization is too strong) or large (the NF has a stronger ability to learn high-frequency features and has weaker implicit regularization). For the TV regularization of PG (see Fig. S3 and Sec. S3.2 in the supplement), a larger $\lambda$ encourages smoother SOS results and reaches optimal performance at a weight of $1\times 10^{-4}$.
@@ 5 Discussion and Conclusion
We have introduced a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and speed of sound in PACT. This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multi-channel image representation analogous to a focal stack. We highlight the following as the key features of our novel method:
@@
Scalability Our efficient combination of coordinate-based SOS representations with a differentiable forward model offers a promising solution for large photoacoustic datasets. By reducing the degrees of freedom and lowering computational costs, the proposed method could enable practical large-scale PACT applications.
Acknowledgments
We gratefully acknowledge the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513. We would also like to thank Liujie Gu and Yan Luo for helping us with the numerical simulations and Yi-Chun Hung and Marcos Ferreira for suggestions on the manuscript.
@@ Acknowledgments
S1.3 In vivo Mouse Liver
The PA signals in our in vivo experiments were collected from a nude mouse liver at 1064 nm [3]. The water temperature was 31C during the experiment, thus we used a background SOS of 1511.4 m/s. See Fig. S2 to see the alignment between recovered SOS and IP, which matches better in our method than others.
Figure S1: (a) Experimentally measured transducer EIR and its Fourier transform. (b) The system MTF is the Fourier transform of the PSF, which is the projection of EIR into image space.
S1.4 Transducer EIR and System MTF
We measured the transducer electrical impulse response (EIR) in the setup we used to collect the data of the leaf phantom and the in vivo mouse liver. We show that our experimentally collected transducer EIR and its Fourier transform is a band-pass filter, shown in Fig. S1a . We calculate the system PSF by projecting this EIR from all 512 transducers into the image space using DAS and calculate the system modulation transfer function (MTF) by taking the 2D Fourier transform of this PSF, shown in  Fig. S1b. The ring structure in the MTF is caused by the projection arcs of DAS. We apply a phase only EIR deconvolution on the collected PA signals before feeding them into our algorithm as a calibration in the phantom and the in vivo experiment.
@@
S1.5 Image Stitching
@@
One critical contribution of our work is the improved wavefront model compared to the previous work APACT [5]. As explained in the main text, the wavefront model represents the wavefront advancement or retardation induced by a nonuniform speed of sound in tissue relative to an ideal spherical wave in a homogeneous medium. Mathematically, the wavefront function $w(\theta)$, which depicts the propagation length difference relative to the spherical wave when received by the transducers, is a function of the propagation direction $\theta$. This wavefront function can be expanded using a Fourier series:
@@
In Fig. S3, we show the effects of TV regularization weight $\lambda$ with our pixel grid SOS representation. We observe that the IP image quality drops when $\lambda>1\times 10^{-4}$, while the SOS accuracy reaches optimum around $\lambda=1\times 10^{-4}$. Therefore, we select $\lambda=1\times 10^{-4}$ (marked with a green star in Fig. S3) for all reconstructions in our implementation of PG.
@@
We provide six videos showing the convergence process of the IP image and SOS for our methods with our numerical phantom (nf_numerical_phantom.mp4 and pg_numerical_phantom.mp4), leaf phantom (nf_leaf_phantom.mp4 and pg_leaf_phantom.mp4), and in vivo mouse liver (nf_in_vivo_mouse_liver.mp4 and pg_in_vivo_mouse_liver.mp4).
Figure S5: All five numerical phantoms and results. (a) The SOS maps used in our numerical phantoms. (b) Reconstructions of our method with NF. (c) Reconstructions of our method with PG. (d) Reconstructions of APACT [5].